ID |
Date |
Author |
Type |
Category |
Subject |
3401
|
Wed Aug 11 16:13:59 2010 |
Jenna | Update | elog | elog restarted | The elog crashed, so we restarted it again. |
3400
|
Wed Aug 11 15:27:16 2010 |
Jenna | Update | Electronics | Rubidium clock phase noise | We unsynched the clocks by unhooking the 1pps locking. I've added it to the plot of the other measurements here, and we've divided by a factor of sqrt(2) in the calibration to get the phase noise from just one clock, so the calibration is now
pi (rad) /6415 (counts) /100/sqrt(2).
I've also added the noise of the clock according to SRS to the plot.
The units of this plot are rad/rt(Hz). I've no idea why it just says magnitude. |
Attachment 1: 08_11_Rb_withspec.pdf
|
|
3399
|
Wed Aug 11 13:28:44 2010 |
josephb | Update | CDS | New cdsIPCx parts, old part breaks models | While working on a test stand at LLO, I've discovered that there's a new update to the Real Time Code Generator that breaks virtually all of our current models.
Previously, we've been using the cdsIPCx part as a flexible link between models, and could be set to RFM (reflected memory), SHMEM (shared memory on the local computer), or PCIE (pci express or dolphin I think) depending on what was written in the IPC file (found in /cvs/cds/caltech/chans/ipc/C1.ipc).
Recently, three new parts were added called cdsIPCx_SHMEM, cdsIPCx_RFM, cdsIPCx_PCIE, and the previous cdsIPCx broken. The cdsIPCx_***** has to match the type written in the IPC file or else it errors out with a rather unhelpful error message which doesn't tell you which part/line is in conflict...
i.e.
***ERROR: IPCx type mis-match: I vs. ISHM
make: *** [g1lsc] Error 255
In anycase, when using a current checkout (updated or checked out after ~July 30th), all cdsIPCx blocks in the simulink models needs to be changed to the approriate type. I'm trying to get a new CDS_PARTS.mdl file I updated into the svn, but for now you can just open up the model file directly in the /advLigoRTS/src/epics/simLink/lib directory and copy it in from there (i.e. cdsIPCx_SHMEM.mdl) to your model file. I'm also trying to get the feCodeGen.pl file changed so the error message actually tells you what channel/part is mismatching so you don't have to guess.
An easy way to modify a file for all shared memory connections without doing a ton of cutting, pasting and renaming is:
sed -i 's/"cdsIPCx"/"cdsIPCx_SHMEM"/g' file_name_here.mdl |
3398
|
Wed Aug 11 12:58:56 2010 |
Jenna | Update | Electronics | Rubidium clock phase noise | I took some measurements of the clock this morning, first without the box, then with the box, then without the box again. All the noise levels look pretty much the same. When I first put the box on, it was only propped up on one side, so I think the clocks got a bit overheated and the data looks ridiculous, which is the first plot. I took it off and let them cool down a bit, and then put the box on, this time with a generous 3 inch gap at the bottom all the way around, and it seemed to be fine after that.
The calibration for the data is pi (rad) /6415 (counts) /100.
Aidan: I edited this post to change the plots from Postscripts to PDFs. |
Attachment 1: 08_11_Rb_crazybox.pdf
|
|
Attachment 2: 08_11_Rb_comparison.pdf
|
|
3397
|
Wed Aug 11 11:51:45 2010 |
Gopal | Update | WIKI-40M Update | 8.5.10 - 8.11.10 Weekly Update | Summary of this Week's Activities:
Thursday, August 5:
X-Displacement Transfer Function Measurement
JPL Tour
Friday, August 6:
Y-Displacement Transfer Function Measurement
Z-Displacement Transfer Function Measurement
Monday, August 9:
Worked on COMSOL/MatLab Interface --> problems may be due to older version
Discussed with Koji options for calling our COMSOL sales representative
Jan and I decided that there is in fact something wrong with the installations on both my Mac and Kallo
Reinstalled on both machines, but the problem was not solved
Jan said we'd go see Larry tomorrow
Tuesday, August 10:
Attempted to figure out Time-Dependent Modal Analysis --> don't think it's what we need
Began reading the LiveLink for MatLab documentation --> even the directions in this produced issues
Discovered "Prescribed acceleration" option for gravity:
A test with it on the simplest stack eliminated the unwanted oscillation, which I guess is a partial success... 
Trying the same thing with Koji on a simple pendulum, however, didn't produce the expected increase in resonant frequency
(Jan was unable to see Larry today, but we're meeting on Wednesday instead).
Wednesday, August 11 (morning):
Some background research on multiple-layer stack theory
Began working on presentations
|
3396
|
Wed Aug 11 02:44:37 2010 |
Zach | Update | elog | elog restarted | You'll never guess what happened: the elog crashed!
I restarted it with the script. Yay. |
3395
|
Tue Aug 10 22:40:55 2010 |
Koji | Summary | PEM | Accelerometer located on and below the PSL table | Result of the accelerometer measurement
Introduction
We wanted to characterize the PSL table before the work before its lifting up.
We put a set of three-axis Wilcoxon accelerometers on the ground and another set on the PSL table through the weekend.
Result
- The data at 9th Aug 00:00(UTC) is used. This was Sunday 5PM in the local time.
- The freq resolution was 0.01Hz. The # of avg was 50.
- The accelerometer signals were calibrated by the value 1.2e-7 V/(m/s^2). We use this absolute value of the spectrum for the comparison purpose.
- The accelerometers were aligned to North(X), East(Y), and Up(Z). There was the coherence observed from 2~20Hz.
The transfer functions are valid only this frequency region although we still can set the lower bound of them.
- The transfer functions in the horizontal directions show huge peaks at around 20Hz. The Q of the peaks are ~30 to ~100.
The vertical transfer function shows somewhat lower peak at around 50Hz with Q of ~10.
Some thoughts
- The low resonant freq and the high Q of the horizontal mode comes from the heaviness of the table.
- We are going to raise the table. This will usually mean that we get the lower resonant freq. This is not nice.
- So, the decision to use 6 tripods rather than 4 was right.
- The steel tripods are expected to give both more rigidity and more damping than the chep-looking hollow Newport legs.
- Concrete grouting of the tripods will also lower the effective height and will benefit for us.
|
Attachment 1: PEM_100809.pdf
|
|
3394
|
Tue Aug 10 17:21:48 2010 |
steve | Update | General | PSL optical table is at 30" height | Mike Gerfen, Koji and Steve,
Our old ( 6 x 10 x 1 ) ft Newport optical table will have the same height as the neighboring AP table when this job is finished.
Today's work was to elevate the optical table ~ 2400 lbs to 30" clearance to create a workable space below.
Atm1, enclosure frame was raised by 6"
Atm2, 2 pellet jacks and 4x4 technology in the work.
Atm3, ready for drill and tap with jig plate in place, note 4 Al legs 8" OD. 1/4" wall safely supporting the load
Atm4, showing clamps that holding 1/2" plywood with jigs
Atm5, We removed the 4 old legs: 6" OD steel, 1/4" wall and 10" high after they were unloaded. |
Attachment 1: P1060586.JPG
|
|
Attachment 2: P1060583.JPG
|
|
Attachment 3: P1060591.JPG
|
|
Attachment 4: P1060593.JPG
|
|
Attachment 5: P1060597.JPG
|
|
3393
|
Tue Aug 10 16:55:38 2010 |
Jenna | Update | Electronics | c1iovme restarted | Alastair and I restarted the c1iovme around the time of my last elog entry (~3:20). |
3392
|
Tue Aug 10 15:23:35 2010 |
Jenna | Update | Electronics | Rubidium clock time constant | [Jenna & Alastair]
We changed the locking time constant on one of the Rubidium clocks using the RbMon software that came with it. We had to use the ancient Dell laptop latitudeD810 because it has a serial port built in, and we couldn't get the usb->serial adapter to work right with the clock. We tried the usb connector on more than one computer, and we had installed the right adapter and the computer seemed to recognize it fine, it just wouldn't communicate with the clock. We even tried it with the Dell latitute laptop and it still failed to work, so the only way seems to be to use the serial port directly.
The clock has a default time constant of 18.2 hours because it's designed to be locked to a GPS clock which is less stable than the Rb clock itself, so we changed it to a time constant of .57 hours. We also changed the length of the BNC cables to get the DC offset to 10mV, but then as I was typing this, we opened up data viewer to look at the real time data and saw the output suddenly leap up, and found that the offset is now -5mV mysteriously, so we went to investigate and found that the gain of the SR560 was still set to 1 from a calibration. We beat one of the clocks with a marconi for a few minutes with the gain still at this level to do another calibration, and then hooked the clocks back up together and upped the gain to 100. The DC offset is currently about 2.5mV. We're going to leave them alone for a few hours, and then check to see what the signal looks like over that period. |
3391
|
Tue Aug 10 05:56:07 2010 |
Koji | Update | PSL | PSL Table Lifting Preparation | Work on Aug 9th
Steve, Jenne, Koji, Alberto, Aidan, Jan, Sharmila, Katherine
From 9am to 6pm
- Shutting down the PSL after the 90885 hours of service
- Removals
- The accelerometers
- The reference cavity chamber
- The periscopes were left so far
- The cables between the PSL table and the outside have been disconnected
- Listed items on Friday
- Some unlisted items recorded and disconnected
- Drained the cooling water from the chiller lines
- Pulled out the chiller connections at the control room as well as the chiller control cables (temp sens & RS232C)
- Sealed the PSL table with plastic sheets
- Put antistatic films to the table (they are supported by the long optical poles)
- Used capton tapes to fix the films on to the table
- Put the white huge plastic sheet to cover the table at once
- Some spaces at the edge of the tables are left flat such that the C-clamps can be attached
- Sealed the AP and the ITMY tables by capton tapes
Some photos are attached in this entry. All of the photos found in the picasa album (click the slideshow)
|
Attachment 1: IMG_2686.jpg
|
|
Attachment 2: IMG_2691.jpg
|
|
Attachment 3: IMG_2692.jpg
|
|
Attachment 4: IMG_2711.jpg
|
|
3390
|
Tue Aug 10 01:05:17 2010 |
Zach | Update | elog | elog restarted | Elog was down. Restarted with the script. |
3389
|
Mon Aug 9 21:50:50 2010 |
nancy | Update | IOO | Mode Cleaner ASC |
Quote: |
The WFS and QPD servos were working. That was great.
Everything was fine except for the time series plots.
I could not get what story you are telling with the time series.
(e.g. your's are good or bad or anything)
|
Well, the data is kind of not enough to be analysed in time domain,
But by far from what I analyse, I think that the new control is not worse than the old one.
I donot also find any better results, except for this one being theoritically stronger. |
3388
|
Mon Aug 9 15:54:43 2010 |
Koji | Update | IOO | Mode Cleaner ASC | The WFS and QPD servos were working. That was great.
Everything was fine except for the time series plots.
I could not get what story you are telling with the time series.
(e.g. your's are good or bad or anything) |
3387
|
Mon Aug 9 13:32:02 2010 |
nancy | Update | IOO | Mode Cleaner ASC | E-log entry for Friday - will attach more plots to this entry on wednesday after i am back to the 40.
Started working at some 1030 hrs and recording the Open Loop Tfs for all 6 loops.
The control was not so good, and I lost the lock quite a number of times while measureing
WFS did not converge when the spot was aligned to the center. But there was convergence to a non-center point. So if the control system was switched on near those points, it was converging to that point.
Autolocker : switches WFS control on directly, whereas the best way is to gradually increase the gain to 1. Also, the autolocker code now needs to be changed to incoporate the switing off the MC2 oplev in down and switch it on in the up script.
After Koji locked the Reference Cavity in the evening, I resumed measurements for the Open Loop TFs.
Measurement of the Open Loop Transfer Functions :
noise waveform was generated using arbitrary wf generator and injected into each loop.
An LPF was applied to have max co-relation at minimum disturabnce. (thanks to Rana)
The Transfer functions, Co-relations and Power Spectra were then measured using the DTT.
Power Spectrum of the IN1, IN2 and EXC shows clearly the suppression of the noise, and OLTF shows the phase margins.
- Courtesy Rana again for suggesting the idea of plotting power spectra of all signals in the same graph.
Later in the night , Koji worked with me and we reflected upon all TFs and changed gains whereevr required according to the phase margin considerations from the Open Loop TFs.
We used the same output matrix given in the previous e-log.
Final gains -
Alignment Gain in the WFS Master - 1.000
Loop |
Gain |
WFS1 P |
0.27 |
WFS1 Y |
0.7 |
WFS2 P |
0.15 |
WFS2 Y |
0.110 |
MC 2OPLEV P |
-0.1 |
MC2 OPLEV Y |
-0.1 |
this measurement invloved locking the MC to the correct position, with the spot centered at both the WFS and the QPD. invloved some cheating (offsets) after we tried centering w/o offsets.
demod signal was also centered while alignment.
credits to Koji for getting the correct lock position and also staying with me till late night in the lab
Important Points to be noted
1. All loops' histories have to be cleared while swtiching them on.
2. turn the loop output before the loop input so that there is no remnant history in the loop.
2. Alignment gain was gradually increased to 1. and tehn the oplevs turned on.
Later measured teh PSD of 6 error signals under 3 conditions -
New Control ON
New Control OFF
Old Control ON
Also measured the time series for the MC_trans and MC_refl for the 3 conditions.
|
Attachment 5: refl_off.pdf
|
|
3386
|
Mon Aug 9 12:46:24 2010 |
Nancy | Update | IOO | Mode Cleaner WFS |
Yesterday , I put in the Output Matrix, and changed the gain sliders for the 4 WFS loops.
|
From how much to how much have you chnged the gain?
I changed the gains from all 4 0.01 to o.27, 0.23, 0.32 and 0.11 and the main alignment gain to be 0.8
Next we stepped to putting in the gains for the MC2 oplev servo.
|
I like to put the credit to Aidan for teaching Nancy how to use FOTON.
Yes, I am sorry for not mentioning this.
Thanks Aidan
|
3385
|
Sat Aug 7 21:57:56 2010 |
rana | Summary | DMF | seisBLRMS restarted | The green xterm on op540m which is running the seisBLRMS DMF got stuck somehow ~3 days ago and lost its NDS connection. I closed the matlab session and restarted it. Seismic trends are now back online. |
Attachment 1: Untitled.png
|
|
3384
|
Sat Aug 7 21:09:59 2010 |
Dmass | Configuration | Computer Scripts / Programs | eLog changes | I made some changes to the elog on nodus:
- Made a backup of /cvs/cds/caltech/elog/elog-2.7.5/elogd.cfg called elogd.cfg.bk.20100407 in the same directory
- Added a folder: /cvs/cds/caltech/elog/elog-2.7.5/logbooks/EAGER_Lab
- Restarted the elog daemon via the start-elog-nodus script in the elog-2.7.5 directory
I saw that the current version of the elog seems to be in the svn, so tried to svn the changes from nodus via ssh, but got this message:
"svn: This client is too old to work with working copy '/cvs/cds/caltech/elog/elog-2.7.5'; please get a newer Subversion client."
I feel I should svn this but don't want to *&#@ the svn/elog up.
For now I will leave it alone and ask a question: Is the folder /cvs/cds/caltech/elog/elog-2.7.5/ under SVN control? Is it also under CVS control?
TL;DR: New tab added to elog.
|
3383
|
Sat Aug 7 11:07:44 2010 |
Koji | Configuration | PSL | MC kept locked / PMC control gain reduced to +13dB | Jenne asked us to keep th MC locked and let the seismometers happy through this weekend.
Note that the work at the control room and the desks are no problem as far as you are quiet.
Nancy told Jenne and me that she finished the work and reverted the WFS to the old state at 4:30AM.
She could not make the elog as it has been crashed.
MC and old MC WFS looks working as usual.
From 6:40AM to 9:40AM the oscillation of the PMC looks present.
At 10:30AM I reduced the gain of the PMC from +15dB to +13dB. |
3382
|
Sat Aug 7 10:47:38 2010 |
Koji | Summary | elog | elog restarted / source of the trouble eliminated | Nancy notified me that the elog crashed. It was fixed.
I restarted elog, but it kept crashing. Some of the entries on Aug 6th caused the trouble.
I tried to refresh the pictures in entries 3376, 3377, 3378. Still it kept crashing.
I started to dig into the elog file itself. (/cvs/cds/caltech/elog/elog-2.7.5/logbooks/40m/100806a.log )
FInally I found that there was some invalid reply links in the entry 3379.
Date: Fri, 06 Aug 2010 19:29:59 -0700
Reply to: 3379
In reply to: 3379
The entry is refering this entry itself. That is weird. So I deleted the reply-to and in-reply-to lines.
Then elogd got happy.
In fact, 3379 was a dupulication of 3380, so I deleted this entry. |
3381
|
Fri Aug 6 20:00:03 2010 |
Koji | Update | PSL | PSL preparation work | PSL preparation work report
Aug 6th 10am-5pm: (Steve, Jenne, Alberto, Kiwamu, Koji)
- We labeled the cables to be disconnected
- These will be disconnected in order to isolate the PSL table and the frame (housing) from the other part of the lab.
- Upon the labeling we made the list and the map of the cables to be removed.
- On Monday we disconnect those cables one by one accoding to the list.
- The following stuffs have been removed from the PSL table
- The big nuts at the side of the PSL table
- The cable ties on the PSL frame
- Innolight 2W
- The green pickoff optics at the edge
- The optics on the shelf
- The oscilloscopes on the shelf
- The OMC power supply cable was visited.
- The connections to the power supply were removed. There are two HV outputs.
- We put thick and long optical poles
- They are placed at the edge of the table so that we can put the plastic sheets on the table without touching the optics.
Plan on Monday
Aug 9th 9am-5pm: (Steve, Jenne, Alberto, Koji)
- Disconnecting the cables ==> All
- Shutting down the PSL ==> Steve/Koji
- Draining the cooling water ==> Steve
- Removals
- The accelerometers ==> Jenne
- The reference cavity chamber ==> Steve
- The small periscope ==> Alberto
- Sealing of the table with the plastic sheets
- The chiller is planned to go to MIT
|
Attachment 1: PSL_cable_map.pdf
|
|
3380
|
Fri Aug 6 19:46:59 2010 |
Gopal | Update | Optic Stacks | (Much Better Looking) Displacement-Displacment Transfer Functions |
Quote: |
I reran the FDA in COMSOL on the MC1/MC3 Stack and produced the following Displacement-Displacement Transfer Functions:
X-Translational Drive has a blue background
Y-Translational Drive has a red background
Z-Translational Drive has a green background
Obtaining the Displacement-to-Phase part of the Transfer Function still produces difficulties -- I'm still working on the COMSOL-Matlab interface to perhaps better facilitate this.
RA: I have deleted those plots because they weren't transfer functions. Transfer functions must always be the ratio of something to something. For example: if I had a nickel for every bad plot I see, I would be a millionaire. In that example, the transfer function would have the units of nickels/plots. For the stacks, it should be meters/meter.
|
My apologies for the mislabeled axes on my previous plots. They have been corrected to a ratio (in./in.), as Rana so kindly suggested in his helpful, not-at-all-condescending response.
I have chosen to stay in the English system because all of the original stack drawings are in inches as well. |
3378
|
Fri Aug 6 17:47:36 2010 |
steve | Summary | General | cranes load tested at 1998 lbs |
Quote: |
Quote: |
Quote: |
The guy from KroneCrane (sp?) came today and started the crane inspection on the X End Crane. There were issues with our crane so he's going to resume on Monday. We turned off the MOPA fur the duration of the inspection.
- None of our cranes have oil in the gearbox and it seems that they never did since they have never been maintained. Sloppy installation job. The crane oiling guy is going to come in on Monday.
- They tried to test the X-End crane with 2500 lbs. (its a 1 ton crane). This tripped the thermal overload on the crane as intended with this test. Unfortunately, the thermal overload switch disabled the 'goes down' circuit instead of the 'goes up' circuit as it should. We double checked the wiring diagram to confirm our hypothesis. Seems the X-End crane was wired up incorrectly in the first place 16 years ago. We'll have to get this fixed.
The plan is that they will bring enough weight to test it at slightly over the rating (1 Ton + 10 %) and we'll retry the certification after the oiling on Monday.
|
The south end crane has one more flaw. The wall cantilever is imbalanced: meaning it wants to rotate south ward, because its axis is off.
This effects the rope winding on the drum as it is shown on Atm2
Atm1 is showing Jay Swar of KoneCrane and the two 1250 lbs load that was used for the test. Overloading the crane at 125% is general practice at load testing.
It was good to see that the load brakes were working well at 2500 lbs. Finally we found a good service company! and thanks for Rana and Alberto
for coming in on Saturday.
|
Jeff Stinson, technician of KoneCrane inspected the south end crane hoist gear box. This was the one that was really low on oil. The full condition require
~ 950cc of EPX-7 (50-70W) high viscosity gear oil. The remaining 120 cc oil was drained and the gear box cover was removed. See Atm 1
He found the gear box, load brake and gearing in good condition. The slow periodic sound of the drive was explained by the split bearings at Atm 3
The Vertex and the east end crane gear boxes needed only 60 cc oil to be added to each Atm 4 and their drives were tested.
Conclusion: all 3 gear boxes and drives are in good working condition.
Tomorrow's plan: load test at 1 ton and correct-check 3 phase wiring.
|
Atm1, service report: load test were performed at max horizontal reach with 1998 lbs ( American Ton is 2000 lbs)
Vertical drives and brakes worked well. The 5 minutes sagging test showed less than 1 mm movement .
The wiring is correct. Earlier hypothesis regarding the wiring ignored the mechanical brake action.
Our cranes are certified now. Operator training and SOP is in the work.
Vertex Folding I -beam will get latch-lock and the south end I-beam will be leveled.
Atm2, south end
Atm3, east end
Atm4, folding crane at ITMX at 14 ft horizontal reach |
Attachment 1: 0.9ton.PDF
|
|
Attachment 2: P1060523.JPG
|
|
Attachment 3: P1060532.JPG
|
|
Attachment 4: P1060541.JPG
|
|
3377
|
Fri Aug 6 16:20:08 2010 |
rana | Update | General | clocks still not locked fully - phase shifted | I tried to get the clocks to be closer to 90 deg for the relative phase by adding some cable length to one of the lines, but they are still wandering too much. We need to use the serial interface and up the gain in their 1PPS locking loops. |
Attachment 1: Untitled.pdf
|
|
3376
|
Fri Aug 6 15:50:29 2010 |
Gopal | Update | Optic Stacks | (Much Better Looking) Displacement-Displacment Transfer Functions | I reran the FDA in COMSOL on the MC1/MC3 Stack and produced the following Displacement-Displacement Transfer Functions:
X-Translational Drive has a blue background
Y-Translational Drive has a red background
Z-Translational Drive has a green background
Obtaining the Displacement-to-Phase part of the Transfer Function still produces difficulties -- I'm still working on the COMSOL-Matlab interface to perhaps better facilitate this.
RA: I have deleted those plots because they weren't transfer functions. Transfer functions must always be the ratio of something to something. For example: if I had a nickel for every bad plot I see, I would be a millionaire. In that example, the transfer function would have the units of nickels/plots. For the stacks, it should be meters/meter.
|
Attachment 1: MC1_MC3_XTrans.png
|
|
Attachment 2: MC1_MC3_YTrans.png
|
|
Attachment 3: MC1_MC3_ZTrans.png
|
|
3375
|
Fri Aug 6 12:44:29 2010 |
Koji | Update | IOO | MC data taking over the weekend | Question:
Do you like to keep the WFS turned on?
This may change the transfer functions between the ground motion to the angle time by time,
and thus change the TF between the GND and the MCL.
Quote: |
Nancy has the Mode Cleaner for her work for the night, and is going to leave the MC happy, locked, autolocker on, WFS enabled, the works, and write down in the elog the time that she's finished. After that, I'm taking MC/seismic data all weekend long. During the weekend, if at all possible, please don't go into the IFO room, especially near the Mode Cleaner. If you do need to go into the IFO room, please elog the time you went in, and the time you left so I can correlate it with my data. This is actually important, so please stick a quick elog entry in if you even think about opening the doors to the IFO room. It is much appreciated.
|
|
3374
|
Fri Aug 6 12:33:10 2010 |
Koji | Update | IOO | Mode Cleaner WFS |
Yesterday , I put in the Output Matrix, and changed the gain sliders for the 4 WFS loops.
|
From how much to how much have you chnged the gain?
Next we stepped to putting in the gains for the MC2 oplev servo.
|
I like to put the credit to Aidan for teaching Nancy how to use FOTON.
This complete 6 * 6 model does work, and was able to keep the transmission held.
|
This should be in this size:
This complete 6 * 6 model does work,
and was able to keep the transmission held.
Yeeeeah!
|
3373
|
Fri Aug 6 12:22:04 2010 |
Jenne | Update | IOO | MC data taking over the weekend | Nancy has the Mode Cleaner for her work for the night, and is going to leave the MC happy, locked, autolocker on, WFS enabled, the works, and write down in the elog the time that she's finished. After that, I'm taking MC/seismic data all weekend long. During the weekend, if at all possible, please don't go into the IFO room, especially near the Mode Cleaner. If you do need to go into the IFO room, please elog the time you went in, and the time you left so I can correlate it with my data. This is actually important, so please stick a quick elog entry in if you even think about opening the doors to the IFO room. It is much appreciated. |
3372
|
Fri Aug 6 10:44:55 2010 |
nancy, koji | Update | IOO | Mode Cleaner WFS |
Quote: |
Upon Nancy's request, I checked the status of the suspensions.
I found that the power strip of the 1Y4 rack was turned off.
Since it has a over current breaker, I don't know whether it happened by someone or over current.
Anyway, I restarted the sus computers, and now the suspensions are damping as usual.
The MC has been aligned, the auto locker is also working.
Incidentally, I found that the WFS servos are not working. Actually since the last night
It repeated losing lock and unlock.
Probably some values of the matrix or the gain is wrong.
I left the WFS as it is because Nancy will put new values this afternoon.
I will ask her to confirm that the old values work at the end of her work.
|
Yesterday , I put in the Output Matrix, and changed the gain sliders for the 4 WFS loops.
It worked and was keeping the lock for the MC.
I then tested whether the MC1 and 3 were following any change in MC2 alignment. It was indeed workinng,
Next we stepped to putting in the gains for the MC2 oplev servo.
the signs are decided on the basis of convergence, and the magnitude is kept very low, to have a very slow control for MC2.
This complete 6 * 6 model does work, and was able to keep the transmission held.
I also tried poking each mirror in pitcg and yaw, and the cavity comes back to high resonance after some time.
This time is indeed large if the poking is made for MC2, and the transmission comes back to normal after big oscillations.
I tried to measure the Open loop TFs for all these loops yesterday, but somehow could not find a correct excitation.
I will do it today.
Plan ahead :
1. Center the spot on MC2 and the QPD
2. Optimize the gains by looking at response to noise.
3. Measure Power Spectrum Density of each error signal.
|
3371
|
Fri Aug 6 08:09:15 2010 |
steve | Update | PSL | PSL preparation work |
Quote: |
PSL preparation work
Aug 5th 10am-4pm?: (Kiwamu, Alberto, Koji)
Removing the unused cables around the PSL table and the control room
Aug 6th 10am-4pm?: (Kiwamu (ex. noon-2pm), Alberto, Koji, Jenne ('till noon) )
- Labeling the cables to be disconnected / making the records ==> All
- Removals
- the big nuts at the side of the PSL table ==> Steve
- the cable ties on the PSL frame ==> Easy
- Innolight 2W ==> Kiwamu
- the green pickoff optics at the edge, if necessary ==> Kiwamu talking with Steve
- Optics on the shelf ==> Jenne / Koji
- Oscilloscopes on the shelf ==> Jenne / Koji
- CCD camera connections (optional, as far as not critical for the operation)
- Put poles on the table (for the plastic sheet) ==> Alberto / Koji
Aug 9th 9am-5pm: (Steve, Jenne, Alberto, Koji)
- Disconnecting the cables ==> All
- Shutting down the PSL ==> Steve/Koji
- Draining the cooling water ==> Steve
- Removals
- The accelerometers ==> Jenne
- the PSL chamber ==> Steve
- Periscopes ==> Alberto
- Sealing of the table with the plastic sheets
- The chiller is planned to go to MIT
|
Monday, August 9
We should move the reference cavity too. Will this cavity be pumped while relocated?
Check and insure that attached and cut-free cables of PSL have enough room to tolerate the raising of the enclosure by 6"
I had second thoughts about the power line to the OMC. Koji was right, we should disconnect them from the power supplies.
The PSL enclosure doors on the north side will have to be removed some times to move exiting and entering ports.
|
3370
|
Thu Aug 5 22:36:11 2010 |
Koji | Update | PSL | PSL preparation work | PSL preparation work
Aug 5th 10am-4pm?: (Kiwamu, Alberto, Koji)
Removing the unused cables around the PSL table and the control room
Aug 6th 10am-4pm?: (Kiwamu (ex. noon-2pm), Alberto, Koji, Jenne ('till noon) )
- Labeling the cables to be disconnected / making the records ==> All
- Removals
- the big nuts at the side of the PSL table ==> Steve
- the cable ties on the PSL frame ==> Easy
- Innolight 2W ==> Kiwamu
- the green pickoff optics at the edge, if necessary ==> Kiwamu talking with Steve
- Optics on the shelf ==> Jenne / Koji
- Oscilloscopes on the shelf ==> Jenne / Koji
- CCD camera connections (optional, as far as not critical for the operation)
- Put poles on the table (for the plastic sheet) ==> Alberto / Koji
Aug 9th 9am-5pm: (Steve, Jenne, Alberto, Koji)
- Disconnecting the cables ==> All
- Shutting down the PSL ==> Steve/Koji
- Draining the cooling water ==> Steve
- Removals
- The accelerometers ==> Jenne
- the PSL chamber ==> Steve
- Periscopes ==> Alberto
- Sealing of the table with the plastic sheets
- The chiller is planned to go to MIT
|
3369
|
Thu Aug 5 17:59:23 2010 |
Koji | Update | PSL | Cable removal from the control room | [Alberto, Kiwamu, and Koji]
We removed the BNC cables from the control room.
The work was as hard as the one I had when I swept a 300m tunnel with a vacuum...
If we could remove the video cables, that would be a real epoch.
We found that the cabling behind the AP table is still quite ugly....grurrrh |
Attachment 1: IMG_2684.jpg
|
|
3368
|
Thu Aug 5 16:18:37 2010 |
Alastair | Update | General | clocks still not locked fully - hut removed | The rubidium clocks are still not quite locked together, though it is clear that the beat frequency has dropped a lot since yesterday.
I checked on the clocks and the 1pps sync light is on. The clocks are really hot again though despite the gap left at the bottom of the igloo. The side of the clocks were hot enough to not be touchable.
I made the executive decision that I would remove the hut just now. We can let the clocks lock together and then put the hut back on just before measuring. This way the hut will isolate from temperature fluctuations during the measurement but it won't be running at the hotter equilibrium temperature. I hope that the temperature won't change too much during the measurement if we put the hut back on.
RA: Attachment deleted because it was in Postscript format. Also not allowed here are the Stone Tablet and Cave Painting formats. |
3367
|
Thu Aug 5 13:05:53 2010 |
Koji | Update | IOO | Mode Cleaner WFS | Upon Nancy's request, I checked the status of the suspensions.
I found that the power strip of the 1Y4 rack was turned off.
Since it has a over current breaker, I don't know whether it happened by someone or over current.
Anyway, I restarted the sus computers, and now the suspensions are damping as usual.
The MC has been aligned, the auto locker is also working.
Incidentally, I found that the WFS servos are not working. Actually since the last night
It repeated losing lock and unlock.
Probably some values of the matrix or the gain is wrong.
I left the WFS as it is because Nancy will put new values this afternoon.
I will ask her to confirm that the old values work at the end of her work. |
3366
|
Thu Aug 5 11:48:52 2010 |
nancy | Update | IOO | Mode Cleaner WFS |
Quote: |
I calculated the MC1&3 Vs WFS1&2 Output Matrix today from the above measurements with koji's help.
the matrix can be generated from the m file at /cvs/cds/caltech/users/nancy/Align_Matrix/matrix.m
these values were put in, and the direction of control is sort of confirmed. I tried twiddling with the gains in the loop to get a 4*4 stable control, but could not succeed.
the mode cleaner is back locked now, and WFS matrix as well as gains are reverted to the old values. (1.30 am)
The output Matrices are
Pitch
0.724
|
0.197
|
0
|
0
|
-1.448
|
-0.758
|
Yaw
0.919
|
0.139
|
0
|
0
|
-0.0106
|
-0.1245
|
|
I realised today morning that there was a flaw in my calculations for the yaw matrix.
Correcting the values, and also making teh tables more readable.
I will test these values once our computers are back to working condition.
PITCH
|
WFS1
|
WFS2
|
MC1
|
0.724
|
0.1964
|
MC2
|
0
|
0
|
MC3
|
-1.4436
|
-0.756
|
YAW
|
WFS1
|
WFS2
|
MC1
|
0.0710
|
0.1074
|
MC2
|
0
|
0
|
MC3
|
0.0082
|
-0.962
|
|
3365
|
Thu Aug 5 01:29:39 2010 |
nancy | Update | IOO | Mode Cleaner WFS |
I calculated the MC1&3 Vs WFS1&2 Output Matrix today from the above measurements with koji's help.
the matrix can be generated from the m file at /cvs/cds/caltech/users/nancy/Align_Matrix/matrix.m
these values were put in, and the direction of control is sort of confirmed. I tried twiddling with the gains in the loop to get a 4*4 stable control, but could not succeed.
the mode cleaner is back locked now, and WFS matrix as well as gains are reverted to the old values. (1.30 am)
The output Matrices are
Pitch
0.724
|
0.197
|
0
|
0
|
-1.448
|
-0.758
|
Yaw
0.919
|
0.139
|
0
|
0
|
-0.0106
|
-0.1245
|
|
3364
|
Thu Aug 5 00:17:41 2010 |
Koji | Update | PSL | PSL preparation work | We start the work on the cables at around the PSL table.
Aug 5th 10am-4pm?: (Kiwamu, Alberto, Koji)
- Removal of the unused cables around the PSL table and the control room
- Removal of the cable ties on the PSL frame
- Removal of the big nuts at the side of the PSL table
Aug 6th 10am-4pm?: (Kiwamu, Alberto, Koji, Jenne (~noon) )
- Labeling of the cables
- Planning of the disconnection
Aug 9th 9am-5pm: (Steve, Jenne, Alberto, Koji)
- Shutting down of the PSL
- Disconnection of the cables
- Draining of the cooling water
- Removal of the accelerometers
- Removal of the PSL chamber
- Sealing of the table with the plastic sheets |
3363
|
Wed Aug 4 20:58:22 2010 |
Gopal | Update | WIKI-40M Update | 7.28.10 - 8.4.10 Weekly Update | Summary of this week's activities:
7/28: Finished Y-Translational 4-Stack Analysis
"Tapered Cantilever" COMSOL tutorial
Tried (and failed) isolating gravity from oscillation
7/29: Developed tilt/rotation load combinations for torsional inputs and showed these to work in the model
Tried using Normal Vector mode on top plate to obtain output tilts; worked for the rectangular bar, but not for the full stack 
Talked to Jan about a 1st-order alternative to gravity - requires Weak Form (only found in COMSOL 3.5 right now)
Began Z-Translational 4-Stack Analysis -- Ran Overnight
7/30: Progress Report 1st Draft
Completed Z-Translational 4-Stack Analysis
8/1: Progress Report 2nd Draft
8/2: Progress Report 3rd Draft
Submitted Progress Report
8/3: Finalized Eigenfrequency Analysis for MC1/MC3 Stack 
24 Physical Eigenmodes plotted and recorded, as expected
Should be good enough for the final report --> focus on transfer function analysis for the remainder of the SURF
8/4: Prescribed Displacement Tests on Simple Rectangular Block --> shown to better produce displacement-displacement transfer functions
X-to-X Transfer Function seems much better when plotted
Should now be able to do the Displacement portion of Transfer Function Analysis on MC1/MC3 for Translational Modes
(I apologize that this update is a little late) |
3362
|
Wed Aug 4 20:15:07 2010 |
Jenne | Update | PEM | Yay! Guralps work again! | After much hassle, the Guralp cable from the ADC Out of the breakout box to the ADC is fixed, and everything is plugged in and working again. The seismometers are back in their regular positions at the ends of the MC, ready for some excellent seis/MC combo data.
I solidified the change of putting the Gur2Z channel into a different BNC input on the ADC. The C1ADCU_PEM.ini file has been changed so that what used to be the Ranger's channel is now recognized as Gur2Z.
Also, I changed the same .ini file to reflect Koji's move of ACC_MC1_Z to the old AUDIO_MIC2 channel, so now all 6 Accelerometer channels have the same calibrations again.
Another big change is the change from old-left-handed convention to new-right-handed convention. The seismometers are aligned the same way they always have been (with the North-South markers aligned with the MC), but now the North-South output is plugged into the BNC on the ADC that is associated with Gur*_X, and the East-West output is plugged into the ADC channel associated with Gur*_Y. This is true for both Guralp Seismometers.
So, now we have:
Gur1_X = Gur1_NS = ADC#10
Gur1_Y = Gur1_EW = ADC#11
Gur1_Z = Gur1_Vert = ADC#12
Gur2_X = Gur2_NS = ADC#2
Gur2_Y = Gur2_EW = ADC#3
Gur2_Z = Gur2_Vert = ADC#24
SEIS_Ranger_Y = no longer in the .ini file |
3361
|
Wed Aug 4 19:50:58 2010 |
rana | Configuration | Electronics | Rubidium clocks too hot: hut removed | Alastair found that the foam hut that he and Jan put on top of the Rb clocks to temperature stabilize them was too good of an insulator. The Rb boxes had gotten very hot and became internally unlocked as seen on the front panel.
After we let them cool down with the box off, I turned them back on. After several minutes the 'Locked' light came back on. Some minutes after that the '1PPS Sync' light also came on, indicating that the two had become somewhat synchronized. It really means that the frequencies are kind of close: I think its roughly that f1-f2 < 2 mHz.
I put the yellow box back on and have left it with a small gap on the bottom so that the hot air can get out. Hopefully, this will protect the clocks from the wind, but not cause them to overheat.
The signal going to the DAQ right now is DC-coupled, with a gain of 1. The peak-peak beat signal in this situation is 6300 counts.
My guess is that the clocks will by synchronized by tomorrow afternoon so that we can get the measurement done. Please don't disturb the clocks or the yellow box around them. Try to minimize any activity around that area. |
Attachment 1: beat.png
|
|
3360
|
Wed Aug 4 16:52:59 2010 |
Razib, Aidan | Update | Phase Camera | Sideband power measurement (updated) | Aidan and I made some attempt to measure the power of the sidebands so that we can calculate our expected signal strength.
Our setup looks like the following:

As light from the laser is split into two at BS1, the transmitted beam has higher power as our BS1 is only coated for 1064nm. We get two reflected beams from BS1, one reflected of the front surface and the other from the back surface. We took the stronger back reflected beam to the EOM driven at 40 MHz (also at 25 MHz at a later time). The AOM produced a reference beam with 40 .000 005 MHz offset which we recombined with the sidebands obtained from the EOM. The beat produced is sent off to PDA 10CF connected to 4395A spectrum analyzer.
The plots for 40MHz sidebands and 25 MHz sidebands looks like this:

From the above spectra, at 40 MHz sideband regime:
Power of the carrier @ 40 MHz = -39.72 dBm
Power of the sideband @ 80 MHz = -60.39 dBm

At 25 MHz sideband regime,
Power of the carrier @ 40 MHz = -40.22 dBm
Power of the upper sideband @ 65 MHz = -61.72 dBm
Power of the lower sideband @ 15 MHz = -60.99 dBm
Power Measurement:
We made some necessary power measurement using a PD connected to a voltmeter after the EOM and the AOM when the EOM is driven at 40 MHz:
___________________________________________________________
Dark : 0.025 V
AOM on: 4.10 V (EOM blocked)
EOM : 2.425 V (AOM blocked)
___________________________________________________________
From the earlier calculation (ref: Elog entry July 28) the power that we expect to see at the PD is,
P= A_c ^2 + A_r^2 + A_(-sb)^2+ A_sb ^2 +2* A_r* A_sb * cos ( w_(r,sb) t ) , where A_c= carrier; A_r= reference beam; A_sb=Upper sideband; A_(-sb)= Lower sideband, w_(r,sb) = w_r - w_sb
P = A_c ^2 + A_r^2 + A_(-sb)^2+ A_sb ^2 +2* A_r* A_sb , letting cos (w_(r,sb) go to 1) is order to approximate the maximum signal
So the signal that we expect to see relative to the DC ( i.e A_c ^2 + A_r^2 + A_(-sb)^2+ A_sb ^2, the first four terms of the power equation) is,
Sig = 2* A_r* A_sb / { A_c ^2 + A_r^2 + A_(-sb)^2+ A_sb ^2 },
Since the modulation index is small, the power in the sideband is very small compared to carrier and the reference beam. So we can ignore the sideband power for the signal expression.
So,
Sig = 2* A_r* A_sb / ( A_c ^2 + A_r^2 )
So if we want to maximize this signal w.r.t the reference then,
d (sig)/ d(A_r) = 2 { ( A_c ^2 - A_r^2) *A_sb } / {( A_c^2 + A_r^2)} ^2
Thus, the signal is maximized when,
A_r^2 = A_c^2
We adjusted the AOM to be driven at + 7.7 dBM so that the new power at the AOM matched the EOM power, which is 2.397 in the voltmeter.
So the power at both the AOM and the EOM are:
P_AOM = ( V_AOM - V_dark) / (PD responsitivity * Transimpedance gain)
= ( 2.397 - 0.025 ) / ( 0.45 * 1.5 x 10 ^5 )
= 3.51 x 10 ^ - 5 W
P_EOM = (V_EOM - V _dark) / (PD responsitivity * Transimpedance gain)
= ( 2. 425 - .0.025) / ( 0.45 * 1.5 x 10 ^5 )
= 3.55 x 10^ - 5 W
From the spectra of the 40 MHz sideband above, the ratio of the carrier and the sideband amplitude is: A_c / A_sb = 10.8 .
P_EOM = A_c ^2 + 2 A_sb ^2
Therefore, A_sb = sqrt ( P_EOM / 118.64) = 5.47 x 10^ - 4 V/m
Thus, A_c = 5.908 x 10^ -3 V/m
and A_r = sqrt ( P_AOM) = 5.92 x 10 -3 V/m.
This measurement can be used to calculate the signal to contrast ratio (SCR) that we expect to see:
SCR = 2 A_r * A_sb / ( A_c^2 + A_r^2 ) = 0.09
Our next step is to measure the actual signal to constrast ratio as seen by the camera. Details of that will be posted soon. |
3358
|
Wed Aug 4 12:49:42 2010 |
nancy | Update | IOO | Mode Cleaner WFS |
Quote: |
Yesterday, I started twiddling with the Mode Cleaner at about 2 pm.
So the seismic data should be all good before that.
I was using it till about 3.30 am, and then left for the night with locking it and swithcing on back the WFS control
Today morning, I have started twiddling with it again, at about 10.30 am.
About my work with the mode cleaner :
I am primarily exciting the mirrors in pitch and yaw, and trying to measure the response of the WFS and the MC2 OPLEV wrt the excitation.
This thus involves switching off the WFS control while measurement.
After two more of those measurements today, I will get to finding new values for the Output Matrix of the WFS for controlling MC1 & 3, and also, try giving in control to MC2 alignment using OPLEV signals.
|
TFs after the measurement -
In the order - MC1 , MC2 , MC3 -pitch and yaw.
These plots let us know about how do the wavefront sensor signals actually respond to the mis-alignments in the mirrors.
For legibility, legend has been includded in only one plot in each pdf., its typically the same for all 3 plots.
the actual xml files for this measurement are in the directory /cvs/cds/caltech/users/nancy/Align_Matrix/highpower/spot_center
It was made sure before each measurement that the MC is best aligned, the WFS are turned off, and the spots on all 3 QPDs are centered.
|
Attachment 1: pit.pdf
|
|
Attachment 2: yaw.pdf
|
|
Attachment 3: pit.pdf
|
|
Attachment 4: yaw.pdf
|
|
Attachment 5: pit.pdf
|
|
Attachment 6: yaw.pdf
|
|
3357
|
Wed Aug 4 11:10:28 2010 |
nancy | Update | IOO | Mode Cleaner WFS | Yesterday, I started twiddling with the Mode Cleaner at about 2 pm.
So the seismic data should be all good before that.
I was using it till about 3.30 am, and then left for the night with locking it and swithcing on back the WFS control
Today morning, I have started twiddling with it again, at about 10.30 am.
About my work with the mode cleaner :
I am primarily exciting the mirrors in pitch and yaw, and trying to measure the response of the WFS and the MC2 OPLEV wrt the excitation.
This thus involves switching off the WFS control while measurement.
After two more of those measurements today, I will get to finding new values for the Output Matrix of the WFS for controlling MC1 & 3, and also, try giving in control to MC2 alignment using OPLEV signals.
|
3356
|
Wed Aug 4 03:01:56 2010 |
Jenne | Update | PEM | Catastrophic Cable Failures |
Quote: |
After some cable swapping this morning, I have determined which cable is bad. It's the Gur1 cable between the seismometer and the breakout box. This is a milspec -> 37pin d-sub cable. I'll pull out the cable and have a look at it after lunch.
|
So, I was wrong about which cable it was, probably in my rush to get some lunch. The actual culprit was the octopus cable that Bob made waaay back in the day (~2 years ago) to go from the "ADC out" of the breakout box (37pin Male Dsub) to 9 BNCs. As it turns out, the Gur1Z channel of that cable was broken on both ends!!!
On one end, we have the 37 pin Dsub. The cable used was so thick (way too thick for this application) that it made a super rigid connection between the wires and the connector, and any bending of the cable stressed this connection, despite the strain-relief of the backshell. The Gur1Z connection snapped off when I was gently wiggling the connections to check them out. Also, since the wires were all so thick, they didn't really fit into the hole in the backshell, so 2 or 3 of them were squished.....straight through the insulation so that several channels were shorted together / potentially shorted to ground. This may explain some of the nasty behavior that Rana and I had seen (although I might have forgotten to elog? My bad.) that even with the inputs of the breakout box all terminated, there was high coherence between the channels. Terminated inputs should give random noise, so this was fishy.
On the other end of the cable we have the 9 BNCs. I had finished redoing the 37pin end of the cable, and was 'beeping' it to check it out, when to my dismay I found that the Gur1Z channels (the inside and the outer shield of the BNC connector) were shorted together. I removed these 2 wires from the Dsub connector to confirm that the BNC was at fault. Koji looked at the BNC with me after I chopped it off of the cable. Bad news strikes again. To get the wires to fit in the inner pin of the BNC connector, the cable-maker had cut off several strands of the wire to make it skinnier. It appears that over the years these cut-off strands wiggled their way to touching the outer shield. This appears to be a danger for all of the BNCs on this cable: a little bit of torque (which one might expect during plugging and unplugging a BNC) and the 2 sides of the differential measurement will be shorted together.
I then decided to start afresh and make my own cable. I found some AWG26 8-twisted-pair cable laying around underneath the Yarm (since this was all I could find, I was just going to do the Gur1 and Gur2 channels, and leave out the Gur3's). The 37 Dsub side was easy, but I seem not able to connect such skinny wire to the BNC connectors in a robust way. Since this bad cable has so far cost me ~2.5 16-hour days of grief, I don't want my new version of it to also be bad. At this point, I await the advice of one wiser than I. I think BNC connectors are designed for something a little closer to ~20AWG, but I could be wrong. Also, they are obviously optimized for coax cable. So what I have now is never going to be great. Maybe tomorrow I can go to the Electronics Shop / Store and buy BNC connectors that are meant to be soldered-to. That would be awesome.
Since I currently have no functional cable to go from Breakout box to ADC, the Guralps are unplugged for tonight.
Conclusions for the day / evening: Frank, Alastair and Jenna are mostly absolved of blame, although the traveling to Bridge and opening and closing the box (which usually involves more plugging and unplugging of cables) probably didn't help this cable out too much. Also, Bob definitely owes me a Sugar Napoleon or something.
In other news, since the Gur2Z ADC channel is totally wacked, I have taken over (but not renamed) the Ranger channel for Gur2Z for now. Jan still has the Ranger hostage over in Bridge, so this is okay for now. |
3355
|
Wed Aug 4 01:44:56 2010 |
Koji | Configuration | PEM | Accelerometer located on and below the PSL table | [Jenne and Koji]
In order to characterize the seismic vibration of the PSL table, we put the accelerometers on and below the PSL table.
On the PSL: MC2 Accelerometers (X, Y, Z) - being connected to CH1-3 of the preamps
East side of the PSL table. X, Y, and Z is directed to North, East, and Up.
On the ground: MC1 Accelerometers (X, Y, Z) - being connected to CH4-6 of the preamps
Beneath the West side of the PSL table. X, Y, and Z is co-aligned to the MC2 ACC.
I found that the C1:PEM-ACC_MC1_Z has large noise in the low freq (~1Hz) region. I tracked down the noise source
and found the noise is still present in the down stream even when the CH17 (C1:PEM-ACC_MC1_Z ) of the ADC IF BNC
(@1Y7) was terminated.
I consulted with Jenne and decided to connect this channel to CH14, which is vacant and has name
C1:PEM-AUDIO_MIC2 (16k).
(for the details of the channel configurations, see /cvs/cds/caltech/chans/daq/C1ADCU_PEM.ini )
I first tried magnetometer channels to steal, but they didn't seem reacting (and the connected to the wrong channels).
I am feeling that we should once entirely check the I/F box.
Note that there looks the difference of the gain by x10 between C1:PEM-ACC_MC* channels and C1:PEM-AUDIO_MIC2 .
[From Jenne: The gain difference is because the C1:PEM-ACC_MC* channels have gain=10 in the .ini file, while C1:PEM-AUDIO_MIC2 uses the default gain=1. ] |
Attachment 1: IMG_2675.jpg
|
|
Attachment 2: IMG_2678.jpg
|
|
Attachment 3: IMG_2682.jpg
|
|
3354
|
Tue Aug 3 11:50:16 2010 |
Jenne | Update | PEM | Seismometer Problem Tracked down | After some cable swapping this morning, I have determined which cable is bad. It's the Gur1 cable between the seismometer and the breakout box. This is a milspec -> 37pin d-sub cable. I'll pull out the cable and have a look at it after lunch. |
3353
|
Tue Aug 3 11:17:10 2010 |
kiwamu | Update | CDS | Diagrams for Cables needed for CDS test | Current Wiring Setup for the Suspension Controls

New Wiring Plan for the Suspension Controls with the New CDS

Missing Stuff for the CDS test
Ideally we can reuse the existing cables, but some of them may not be long enough for the new wiring.
The diagram below shows extremely non-ideal case.

Some more information will be summarized on the wiki later.
|
3352
|
Tue Aug 3 03:15:06 2010 |
nancy | Update | IOO | MC back to locked mode | I turned the WFS gain to 0.02 back, and the MC is locked, the data for the seismic motion might be meaningful nowforth. |
3351
|
Tue Aug 3 01:51:01 2010 |
Jenne | Update | PEM | Seismometer Update: Still not good, but perhaps getting better |
Quote: |
Today's seismometer diagnosis activities are still underway, this is just an update (since I did some reboots):
Problem 1: X and Z channels on both seismometers were flipped. I unplugged an X cable (East/West on the cable labels) and the Z channel (vert) would go to floating ADC zero. Rana suggested that the ADCs sometimes have random channel hopping, and that a reboot of the c0dcu1 computer which handles the PEM ADCU should fix this problem. I keyed the c0dcu1 / c0daqawg crate, those computers came back just fine, and the channels were no longer flipped. This is a good thing. Although now it's actually the Z channels that were / are bad on both seismometers, not the X's.
While rebooting those computers, c1iovme, c1sosvme, c1susvme1 and c1susvme2 crashed. I rebooted them, although for the first few power cycles c1susvme1&2 couldn't mount /cvs/cds/caltech. Eventually they did, and life is good again. Except that the seismometers are still funny.
|
Some more progress, but still not complete:
Jan and I looked at all of the Gur channels on a 'scope (battery, so as not to be grounded), and 5 of the 6 looked good. We were looking at the BNCs just as they go into the ADC. The one which still looks bad is Gur1Z. The 'scope just doesn't see any signal on that channel.
In addition, the ADC's BNC input #4 (which normally has Gur2Z) looks totally shot. When it's floating, the signal on dataviewer definitely doesn't look floating. I'm probably going to have to move over to another channel, and just give that one up (this ADC already has several channels which have been declared bad, so maybe it's not a surprise that this can happen?)
Since one of the Gur signals looks bad (Gur1Z) and one of the ADC channels looks bad (usual Gur2Z), I switched the Z channels on the ADC board, so the channel being saved as Gur1Z is in fact Gur2Z. This is valid as of ~1:15am until further elog notice.
During my investigations into why Gur1Z is funny, I also looked at the signal on the BNC octopus cable coming straight from the output of the Guralp Breakout Box (this is the cable which goes from "ADC Out" on the back of the box which is a 37 pin D-sub to 9 differential BNCs), and sometimes I saw zero on the 'scope, but sometimes there was a signal which would coincide with jumping tests. Whenever there was a signal however, it was always a way lower amplitude (at least by a factor of 10?) than the other channels.
All of this craziness led to me pulling the Guralp box to investigate.
Upon opening the box, I recalled that the channels go in order: Vert, NS, EW. The Gur1Z channel is that first vert channel, and it's the one which always had a blue input capacitor rather than a surface mount one. Being suspicious of Frank and Alastair, since they seemed unhappy with my capacitor choices, I wondered if they had wiggled the blue cap, and tore something loose. Just in case, and to make things seem more uniform, I replaced the blue cap with a surface mount 1uF cap. (Actually its 0.909uF, replacing the 0.905uF blue cap, according to the black DMM that measures capacitance.) While I was in there, since it had been a problem in the past (elog 2811), I relflowed the solder on some of the resistors, especially near the output op amp.
Anyhow, none of that may have been necessary. All 6 of the Gur channels were examined on a 'scope, using clip doodles to measure the various Test Points on the circuit. I looked at all of the TPs in Gur1Z, and I didn't find that any particular stage was any noisier than the others. Also, all 6 of the Gur channels seemed totally fine in terms of sending a good signal to the output of the box, including Gur1Z which is currently under investigation. All of the channels passed the "output looks ~20x the input" test, and for approximately equal thumping on the ground all 6 channels seemed to have similar amplitude outputs. The Z channels on both channels one and channels two were a little bigger than the X's or Y's, but the 2 Z channels were about the same. This test was done using Guralp2 and the Gur2 cable on both channels 1 and 2, and then checked with Guralp 1, using the Gur2 cable on channels 1. The Vert1 channel always seemed good.
I now am suspicious of one or more of the cables: either the Gur1 cable from the seismometer to the box, or the Vert1 channel of the octopus cable. I'm satisfied that the BNC cables running through the cable tray are okay (although it might not hurt to check that they all successfully send a sine wave...) I opened up the backshell of the Gur1 cable, on the end that connects to the breakout box. Nothing seemed amiss. I still need to Beep the cable to check its connections, and look at the octopus cable.
Recap / Current Status: Breakout Box is reinstalled, both seismometers hooked up. The Z channels on the seismometers are swapped at the ADC input. The dataviewer channels Gur1_X, Gur1_Y, Gur1_Z (which is actually Z of Gur2 seismometer) and Gur2_X, Gur2_Y are all good. Nancy is going to leave the MC in a happy place, and note the time when she's done. Tomorrow I'll check out the cables for the Gur1Z seismometer channel. |
3350
|
Mon Aug 2 21:52:57 2010 |
Koji | Update | IOO | MC is running at the full power | [Nancy and Koji]
We restored the full power operation of the MC.
Restoration of the suspensions
- Found the suspension watch dogs are left turned off.
- Found c1susvme1/2 were not running.
- Launched the realtime processes on c1susvme1/2 and c1iscey
- Restored the watch dogs. The suspensions looked fine.
Preparations for the high power
- Put an ND2.0 before the MCT CCD. Confirmed the ND reflection is damped.
MCT QPD is not necessary to be touched.
The high power operation of the MC / post lock adjustment
- Locked the MC under the autolocker being disabled.
- Adjusted the aperture on the MC2 face camera
- Adjusted the spot positions on the WFS QPDs
- Reverted the scripts to the high power ones
(mcup / mcdown / autolockMCmain40m)
- Logged in to op340m and restarted autolockMCmain40m
The autolocker seems working correctly. |
|